OpenStudy (dtan5457):

Trig question for tan(a+b) evaluate the expression given that cos a=4/5 with 0

OpenStudy (dtan5457):

for the tan sum formula i'm getting -9/8 for the numerator

OpenStudy (dtan5457):

but the denominator im not sure

OpenStudy (dtan5457):

is it 77/32..?

OpenStudy (dtan5457):

@Nnesha @freckles

OpenStudy (dtan5457):

@sleepyjess

OpenStudy (anonymous):

\[\tan(a+b)=\frac{ \sin(a+b)}{ \cos(a+b) }\]

OpenStudy (anonymous):

\[\sin(a+b)=sina*cosb+sinb*cosa\]

OpenStudy (anonymous):

\[\cos(a+b)=cosa*cosb-sina*sinb\]

OpenStudy (anonymous):

if the sinb is -15/17 when b is in IV quater, then sinb=15/17 when b is in I and II quaters

OpenStudy (dtan5457):

can we use this formula

OpenStudy (dtan5457):

got tanA=3/4 tanb=-15/8

OpenStudy (anonymous):

yes we can, but we are given cos and sin, so i used what i wrote

OpenStudy (dtan5457):

but im plugging it in wrong or something because the answer i got is -36/77 but the answer key was 117/4

OpenStudy (dtan5457):

what are you getting?

OpenStudy (anonymous):

that's true i also got that. I don't know then. sorry maybe these can help @SolomonZelman @TheSmartOne @amistre64 @freckles @perl

OpenStudy (dtan5457):

the answer key is probably wrong

OpenStudy (dtan5457):

I wouldn't be suprised if it is

OpenStudy (anonymous):

maybe.

OpenStudy (dtan5457):

ill stick with my answer

OpenStudy (dtan5457):

ty

OpenStudy (anonymous):

np :)

OpenStudy (freckles):

@dtan5457 what did you get for sin(a) and cos(b)?

OpenStudy (dtan5457):

3/5, 8/17, but i still used the tan formula

OpenStudy (freckles):

ok that is great so tan(a)=sin(a)/cos(a)=3/4 and tan(b)=sin(b)/cos(b)=-15/17

OpenStudy (freckles):

or these the numbers you input?

OpenStudy (freckles):

oops -15/8*

OpenStudy (dtan5457):

yeah

OpenStudy (dtan5457):

final answer i got was -36/77

OpenStudy (dtan5457):

can you confirm?

OpenStudy (freckles):

\[\frac{\frac{3}{4}-\frac{15}{8}}{1-\frac{3}{4} \frac{-15}{8}}=\frac{8(3)-4(15)}{4(8)+3(15)} \\ =\frac{24-60}{32+45}=\frac{-36}{77}\]

OpenStudy (dtan5457):

always knew i was right